# Interplay of Zeeman field, Rashba spin-orbit interaction, and superconductivity: Transition temperature and quasiparticle excitations

### v1

@@ Appendix G Bogoliubov transformation: $p$-wave pairing states with only Rashba SOC
$v_{-\mathbf{k}}^{*}$ $=\frac{1}{2}\begin{pmatrix}-e^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-}+\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+}&e^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+}+\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-}\\ e^{-i\varphi_{\mathbf{k}}}(-e^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-}-\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+})&e^{-i\varphi_{\mathbf{k}}}(e^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+}-\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-})\end{pmatrix},$ (99)
$u_{-\mathbf{k}}^{*}$ $=\frac{e^{-i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}-e^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}-}-\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}&e^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}+}-\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\\ e^{-i\varphi_{\mathbf{k}}}(-e^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}-}+\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+})&e^{-i\varphi_{\mathbf{k}}}(e^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}+}+\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-})\end{pmatrix}.$
Appendix H Continuous phase transitions in $p$-wave pairing states: the superconducting $T_{c}$
@@
H.1 ESP states: The effects of a sole perpendicular Zeeman field
@@
$\Delta=\sum_{\mathbf{k}}\frac{3}{2}V_{1}\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}\{(U_{\mathbf{k}})_{31}^{*}(U_{\mathbf{k}})_{11}f(E_{\mathbf{k}+})+(U_{\mathbf{k}})_{32}^{*}(U_{\mathbf{k}})_{12}f(E_{\mathbf{k}-})+(U_{\mathbf{k}})_{33}^{*}(U_{\mathbf{k}})_{13}[1-f(E_{-\mathbf{k}+})]+(U_{\mathbf{k}})_{34}^{*}(U_{\mathbf{k}})_{14}[1-f(E_{-\mathbf{k}-})]\}.$ (101)
Substituting the Bogoliubov transformation matrices given in Appendix F into the above equation, we have
$\Delta=$ $\sum_{\mathbf{k}}\frac{3V_{1}}{2}\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}\cdot\frac{1}{4}\sum_{s=\pm}\left[\frac{d_{3}(E_{\mathbf{k}\parallel}+s\mu_{B}H_{x})}{E_{\mathbf{k}s}E_{\mathbf{k}\parallel}}+\frac{id_{2}}{E_{\mathbf{k}s}}\right]\tanh{\frac{\beta E_{\mathbf{k}s}}{2}}$ (102)
@@
To obtain the superconducting $T_{c}$ of a continuous phase transition, we take the limit $\Delta\to{}0$ and put $E_{\mathbf{k}\parallel}$ and $E_{\mathbf{k}\pm}$, and $d_{2}$ and $d_{3}$ given in Appendix F into the above gap equation. The resulting $T_{c}$ equation is the following
@@
$\frac{T_{c}}{T_{c0}}$ $=\frac{\pi e^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{\mu_{B}H_{x}},$
@@
$\frac{T_{c}}{T_{c0}}$ $=\sqrt{3}e^{-\frac{5}{6}}\cdot\frac{H_{0}}{H_{x}},$
@@
$k_{B}T_{c0}=\hbar\omega_{D}\left(\frac{2e^{\gamma}}{\pi}\right)e^{-\frac{1}{N(0)V_{1}}}$
@@
$\Delta_{0}=\hbar\omega_{D}{}e^{\frac{5}{6}}e^{-\frac{1}{N(0)V_{1}}},$
@@
H.2 ESP states: The effects of only Rashba SOC
Then we consider the four ESP states with only Rashba SOC. Substituting the Bogoliubov transformation matrices given in Appendix G into the gap equation, we have
@@ H.2 ESP states: The effects of only Rashba SOC
$=$ $\sum_{s=\pm}\frac{e^{i\gamma_{\mathbf{k}}}}{4}\left[\frac{d_{3}\sin\theta_{\mathbf{k}}}{E_{\mathbf{k}s}}-\frac{d_{1}\cos\theta_{\mathbf{k}}}{E_{\mathbf{k}\perp}}\cdot\frac{E_{\mathbf{k}\perp}+sgk_{F}}{E_{\mathbf{k}s}}+\frac{id_{2}}{E_{\mathbf{k}\perp}}\cdot\frac{E_{\mathbf{k}\perp}+sgk_{F}}{E_{\mathbf{k}s}}\right](-1)\tanh\frac{\beta E_{\mathbf{k}s}}{2}.$
@@
$\frac{T_{c}}{T_{c0}}$ $=\left[\frac{\pi e^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{1}{4}},$
@@
(2) For the $k_{y}\hat{x}-k_{x}\hat{y}$ pairing state, the $T_{c}$ equation for a seconder-order phase transition is
@@
$\frac{T_{c}}{T_{c0}}$ $=\left[\frac{\pi e^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{3}{2}},$
@@
H.3 OSP states: The effects of only Rashba SOC
@@
Putting the Bogoliubov transformation matrices in Appendix G into the gap equation leads to
@@
=±$=\sum_{\pm}-\frac{e^{i\eta_{\mathbf{k}}}}{4}\left[\frac{d_{3}}{E_{\mathbf{k}\pm}}\left(\cos\theta_{\mathbf{k}}\pm\frac{\xi_{\mathbf{k}}}{E_{\mathbf{k}\perp}}\right)+\frac{d_{1}\sin{\theta_{\mathbf{k}}}}{E_{\mathbf{k}\perp}}\cdot\frac{E_{\mathbf{k}\perp}\pm gk_{F}}{E_{\mathbf{k}\pm}}\right]\tanh{\frac{\beta E_{\mathbf{k}\pm}}{2}}.$
@@
$\frac{T_{c}}{T_{c0}}=\left[\frac{e^{-\gamma}}{4}\cdot\frac{2\pi k_{B}T_{c0}}{gk_{F}}\right]^{\frac{2}{3}},$

### v2

@@ Appendix H Bogoliubov transformation: $p$-wave pairing states with only Rashba SOC
$v_{-\mathbf{k}}^{*}$ $=\frac{1}{2}\begin{pmatrix}-\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-}+\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+}&\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+}+\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-}\\ \mathrm{e}^{-i\varphi_{\mathbf{k}}}(-\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}-}-\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}+})&\mathrm{e}^{-i\varphi_{\mathbf{k}}}(\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}+}-\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}-})\end{pmatrix},$ (99)
$u_{-\mathbf{k}}^{*}$ $=\frac{\mathrm{e}^{-i\gamma_{\mathbf{k}}}}{2}\begin{pmatrix}-\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}-}-\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+}&\mathrm{e}^{i\Omega_{\mathbf{k}}}\cos{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}+}-\sin{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-}\\ \mathrm{e}^{-i\varphi_{\mathbf{k}}}(-\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}2+}b_{\mathbf{k}-}+\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}1+}b_{\mathbf{k}+})&\mathrm{e}^{-i\varphi_{\mathbf{k}}}(\mathrm{e}^{i\Omega_{\mathbf{k}}}\sin{\frac{\theta_{\mathbf{k}}}{2}}a_{\mathbf{k}1-}b_{\mathbf{k}+}+\cos{\frac{\theta_{\mathbf{k}}}{2}}s_{\mathbf{k}3}a_{\mathbf{k}2-}b_{\mathbf{k}-})\end{pmatrix}.$
Appendix I Continuous phase transitions in $p$-wave pairing states: the superconducting $T_{c}$
@@
I.1 ESP states: The effects of a sole perpendicular Zeeman field
@@
$\Delta=\sum_{\mathbf{k}}\frac{3}{2}V_{1}\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\{(U_{\mathbf{k}})_{31}^{*}(U_{\mathbf{k}})_{11}f(E_{\mathbf{k}+})+(U_{\mathbf{k}})_{32}^{*}(U_{\mathbf{k}})_{12}f(E_{\mathbf{k}-})+(U_{\mathbf{k}})_{33}^{*}(U_{\mathbf{k}})_{13}[1-f(E_{-\mathbf{k}+})]+(U_{\mathbf{k}})_{34}^{*}(U_{\mathbf{k}})_{14}[1-f(E_{-\mathbf{k}-})]\}.$ (101)
Substituting the Bogoliubov transformation matrices given in Appendix G into the above equation, we have
$\Delta=$ $\sum_{\mathbf{k}}\frac{3V_{1}}{2}\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\cdot\frac{1}{4}\sum_{s=\pm}\left[\frac{d_{3}(E_{\mathbf{k}\parallel}+s\mu_{B}H_{x})}{E_{\mathbf{k}s}E_{\mathbf{k}\parallel}}+\frac{id_{2}}{E_{\mathbf{k}s}}\right]\tanh{\frac{\beta E_{\mathbf{k}s}}{2}}$ (102)
@@
To obtain the superconducting $T_{c}$ of a continuous phase transition, we take the limit $\Delta\to{}0$ and put $E_{\mathbf{k}\parallel}$ and $E_{\mathbf{k}\pm}$, and $d_{2}$ and $d_{3}$ given in Appendix G into the above gap equation. The resulting $T_{c}$ equation is the following
@@
$\frac{T_{c}}{T_{c0}}$ $=\frac{\pi\mathrm{e}^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{\mu_{B}H_{x}},$
@@
$\frac{T_{c}}{T_{c0}}$ $=\sqrt{3}\mathrm{e}^{-\frac{5}{6}}\cdot\frac{H_{0}}{H_{x}},$
@@
$k_{B}T_{c0}=\hbar\omega_{D}\left(\frac{2\mathrm{e}^{\gamma}}{\pi}\right)\mathrm{e}^{-\frac{1}{N(0)V_{1}}}$
@@
$\Delta_{0}=\hbar\omega_{D}{}\mathrm{e}^{\frac{5}{6}}\mathrm{e}^{-\frac{1}{N(0)V_{1}}},$
@@
I.2 ESP states: The effects of only Rashba SOC
Then we consider the four ESP states with only Rashba SOC. Substituting the Bogoliubov transformation matrices given in Appendix H into the gap equation, we have
@@ I.2 ESP states: The effects of only Rashba SOC
$=$ $\sum_{s=\pm}\frac{\mathrm{e}^{i\gamma_{\mathbf{k}}}}{4}\left[\frac{d_{3}\sin\theta_{\mathbf{k}}}{E_{\mathbf{k}s}}-\frac{d_{1}\cos\theta_{\mathbf{k}}}{E_{\mathbf{k}\perp}}\cdot\frac{E_{\mathbf{k}\perp}+sgk_{F}}{E_{\mathbf{k}s}}+\frac{id_{2}}{E_{\mathbf{k}\perp}}\cdot\frac{E_{\mathbf{k}\perp}+sgk_{F}}{E_{\mathbf{k}s}}\right](-1)\tanh\frac{\beta E_{\mathbf{k}s}}{2}.$
@@
$\frac{T_{c}}{T_{c0}}$ $=\left[\frac{\pi\mathrm{e}^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{1}{4}},$
@@
(2) For the $k_{y}\hat{x}-k_{x}\hat{y}$ pairing state, the $T_{c}$ equation for a second-order phase transition is
@@
$\frac{T_{c}}{T_{c0}}$ $=\left[\frac{\pi\mathrm{e}^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{3}{2}},$
@@
I.3 OSP states: The effects of only Rashba SOC
@@
Putting the Bogoliubov transformation matrices in Appendix H into the gap equation leads to
@@
=±$=\sum_{\pm}-\frac{\mathrm{e}^{i\eta_{\mathbf{k}}}}{4}\left[\frac{d_{3}}{E_{\mathbf{k}\pm}}\left(\cos\theta_{\mathbf{k}}\pm\frac{\xi_{\mathbf{k}}}{E_{\mathbf{k}\perp}}\right)+\frac{d_{1}\sin{\theta_{\mathbf{k}}}}{E_{\mathbf{k}\perp}}\cdot\frac{E_{\mathbf{k}\perp}\pm gk_{F}}{E_{\mathbf{k}\pm}}\right]\tanh{\frac{\beta E_{\mathbf{k}\pm}}{2}}.$
@@
$\frac{T_{c}}{T_{c0}}=\left[\frac{\mathrm{e}^{-\gamma}}{4}\cdot\frac{2\pi k_{B}T_{c0}}{gk_{F}}\right]^{\frac{2}{3}},$